# Simple is Effective: The Roles of Graphs and Large Language Models in Knowledge-Graph-Based Retrieval-Augmented Generation

### v2

@@ 3.1 Efficient, Flexible and Expressive Subgraph Retrieval
A Lightweight Implementation For $p_{\theta}(\cdot|z_{\tau}({\mathcal{G}},q),q)$ We present a lightweight implementation of $p_{\theta}$ that integrates structural and semantic information. Following previous approaches (Karpukhin et al., 2020; Gao et al., 2024b), we employ off-the-shelf pre-trained text encoders to embed all entities/relations in a KG based on their text attributes. These semantic text embeddings are computed and stored in a vector database during the pre-processing stage for efficient retrieval. For a newly arrived question $q$, we embed $q$ to obtain $z_{q}$ and retrieve embeddings $z_{h},z_{r},z_{t}$ from the vector database for the involved entities and relation. After computing DDEs $z_{\tau}$, an MLP is employed for binary classification using the concatenated input $[z_{q}\|z_{h}\|z_{r}\|z_{t}\|z_{\tau}]$.
@@
Relevant Designs We considered several alternative design options but found them less suitable due to concerns regarding efficiency and adaptability to KG updates. Cross-encoders, which concatenate a question and a retrieval candidate for joint embedding (Wolf et al., 2019), potentially offer better retrieval performance. However, due to the inability to pre-compute embeddings, this approach significantly reduces retrieval efficiency when dealing with a large number of retrieval candidates, as is the case in triple retrieval. Li et al. (2023a) embeds each triple as a whole rather than individual entities and relations. However, this approach incurs higher computational and storage costs and exhibits reduced generalizability to the triples that are new combinations of old entities and relations. Our implementation allows for fast triple scoring while maintaining good generalizability.
@@
Figure 2: The prompt used in SubgraphRAG. Concrete examples can be found in Appendix D).
@@ 3.2 Prompting-Based LLM Reasoning
We utilize an LLM to reason over ${\mathcal{G}}_{q}$ by incorporating a linearized list of triples from ${\mathcal{G}}_{q}$ into the prompt. This enables the LLM to ground its reasoning in the retrieved subgraph and identify the answers $\hat{\mathcal{A}}_{q}$ from the entities within ${\mathcal{G}}_{q}$, addressing issues such as hallucinations and outdated knowledge (Lin et al., 2019; Shuster et al., 2021; Vu et al., 2024). Specifically, we prompt the LLM not only to provide answers but also to generate knowledge-grounded explanations based on the input subgraph. We adopt in-context learning (ICL) (Brown et al., 2020) and design dedicated prompt templates with explanation demonstrations to guide the LLM’s reasoning process (see Fig. 2).
By avoiding the need for fine-tuning LLMs, we reduce computational costs, enable the use of SOTA black-box LLMs, and maintain the framework’s generalizability, even for unseen KGs. While fine-tuning may enhance prediction accuracy, it often diminishes general reasoning and explanatory capabilities. Furthermore, high-quality labels for text-based explanations are typically unavailable in practical question-answering tasks. Consequently, previous KG-based RAG approaches that rely on fine-tuning often generate reasoning explanations using larger, unfine-tuned LLMs, such as GPT-4, to serve as auxiliary labels for additional training (Luo et al., 2024).
Regarding the size $K$ of the retrieved subgraph, while increasing $K$ in principle improves the coverage of relevant information, it also incurs higher costs/latency for LLM reasoning and risks introducing more irrelevant information that may ultimately hurt LLM reasoning (Xu et al., 2024; Liu et al., 2024b). Different LLMs are inherently equipped with different sized context window and also exhibit distinct capabilities in reasoning over long-context retrieval results (Dubey et al., 2024; Leng et al., 2024). As such, although the training of SubgraphRAG retriever is LLM-agnostic, the size $K$ needs to be properly selected per LLM and cost/latency constraint. In Section 4.2, we empirically verify that more powerful LLMs can benefit from incorporating a larger-sized retrieved subgraph, demonstrating the benefit of size-adjustable subgraph retrieval in SubgraphRAG.
3.3 More Related Works
Along with the introduction of components in SubgraphRAG, we have introduced the most relevant works. Other related works are discussed in Appendix A due to the space limitation.
@@ 4 Experiments
Model Triple Recall GPT-4o Triple Recall Answer Entity Recall WebQSP CWQ WebQSP CWQ WebQSP CWQ $1$ $2$ $1$ $2$ $\geq 3$ $1$ $2$ $1$ $2$ $\geq 3$ $1$ $2$ $1$ $2$ $\geq 3$ $(65.8\%)$ $(34.2\%)$ $(28.0\%)$ $(65.9\%)$ $(6.1\%)$ $(65.8\%)$ $(34.2\%)$ $(28.0\%)$ $(65.9\%)$ $(6.1\%)$ $(65.8\%)$ $(34.2\%)$ $(28.0\%)$ $(65.9\%)$ $(6.1\%)$ cosine similarity $0.874$ $0.405$ $0.629$ $0.442$ $0.333$ $0.847$ $0.483$ $0.629$ $0.511$ $0.464$ $0.943$ $0.253$ $0.903$ $0.472$ $0.289$ Retrieve-Rewrite-Answer $0.064$ $0.046$ $0.062$ $0.061$ $0.745$ $0.729$ RoG $0.869$ $0.415$ $0.766$ $0.597$ $0.253$ $0.446$ $0.271$ $0.347$ $0.293$ $0.122$ $0.874$ $0.677$ $0.920$ $0.827$ $0.628$ G-Retriever $0.335$ $0.216$ $0.134$ $0.205$ $0.168$ $0.345$ $0.284$ $0.159$ $0.240$ $0.226$ $0.596$ $0.446$ $0.377$ $0.384$ $0.269$ MLP $0.828$ $0.687$ $0.651$ $0.690$ $0.534$ $0.811$ $0.781$ $0.635$ $0.707$ $0.616$ $0.933$ $0.874$ $0.932$ $0.870$ $\underline{\mathbf{0.793}}$ MLP + topic entity $0.944$ $0.729$ $\underline{\mathbf{0.854}}$ $0.750$ $0.560$ $0.884$ $0.775$ $0.769$ $0.773$ $0.647$ $0.976$ $0.843$ $\underline{\mathbf{0.956}}$ $0.885$ $0.665$ SubgraphRAG $\underline{\mathbf{0.953}}$ $\underline{\mathbf{0.748}}$ $0.831$ $\underline{\mathbf{0.820}}$ $\underline{\mathbf{0.626}}$ $\underline{\mathbf{0.908}}$ $\underline{\mathbf{0.809}}$ $\underline{\mathbf{0.823}}$ $\underline{\mathbf{0.860}}$ $\underline{\mathbf{0.755}}$ $\underline{\mathbf{0.977}}$ $\underline{\mathbf{0.881}}$ $0.946$ $\underline{\mathbf{0.916}}$ $0.741$
@@ 4.1 Evaluation for Retrieval (Q1 & Q2)
Baseline Retrievers. Li et al. (2023a) introduces a structure-free retriever that performs cosine similarity search based on triple embeddings, which we refer to as cosine similarity. Retrieve-Rewrite-Answer (Wu et al., 2023) proposes a constrained path search, predicting relation paths then searching for matched paths. RoG (Luo et al., 2024) adopts a similar strategy but enhances it by fine-tuning an LLM for generative relation path prediction. G-Retriever (He et al., 2024) combines cosine similarity search with combinatorial optimization to construct a connected subgraph.
Implementation Details. We employ gte-large-en-v1.5 (Li et al., 2023c) as the pre-trained text encoder for both the cosine similarity baseline and SubgraphRAG, a 434M model that achieves a good balance between efficiency and English retrieval performance, as evidenced by the Massive Text Embedding Benchmark (MTEB) leaderboard (Muennighoff et al., 2023). For supervision signals, there are no ground-truth relevant subgraphs for a query $q$. Previous path-based subgraph retrievers adopt the shortest paths between the topic and answer entities for the weak supervision signals (Zhang et al., 2022; Luo et al., 2024). We also utilize these shortest paths as the heuristic relevant subgraphs $\tilde{{\mathcal{G}}}_{q}$ to train $\mathbb{Q}_{\theta}$ as in Eq. 2. To reduce the size of candidate triples ${\mathcal{G}}$, we construct subgraphs centered at topic entities ${\mathcal{T}}_{q}$, following previous works. See Appendix B.1 for more details.
@@
WebQSP CWQ Macro-F1 Hit Macro-F1 Hit KD-CoT 52.5 68.6 - 55.7 ToG (GPT-4) - 82.6 - 67.6111Their computation of Hit is different from other baselines and may overestimate the performance. We were unable to reproduce their results following their provided instructions. EtD - 82.5 - 62.0 StructGPT - 74.69 - - Retrieve-Rewrite-Answer - 79.36 - - G-Retriever 53.41 73.46 - - RoG-Joint 70.26 86.67 54.63 61.94 RoG-Sep 66.45 82.19 53.87 60.55 SubgraphRAG + Llama3.1-8B 70.57 86.61 47.16 56.98 SubgraphRAG + Llama3.1-70B 74.70 86.24 51.78 57.89 SubgraphRAG + ChatGPT 69.21 83.11 49.13 56.27 SubgraphRAG + GPT-4o-mini 77.45 90.11 54.13 62.02 SubgraphRAG + GPT-4o 76.46 89.80 59.08 66.69 SubgraphRAG + GPT-4o-mini (200) 77.82 90.54 54.69 63.49 SubgraphRAG + GPT-4o (200) 78.24 90.91 59.42 67.49 SubgraphRAG + GPT-4o-mini (500) 77.67 91.22 55.41 64.97 SubgraphRAG + Llama3.1-8B ($\leftrightarrow$) 66.42 83.42 37.96 48.57 SubgraphRAG + GPT-4o-mini ($\leftrightarrow$) 73.81 88.08 44.69 54.21 SubgraphRAG + GPT-4o-mini ($\leftrightarrow$, 500) 76.20 91.22 50.30 60.80
@@
Evaluation Metrics. Our retrieval evaluation encompasses both effectiveness and efficiency. For effectiveness, we employ three recall metrics: recall of triples in shortest paths (i.e., $\tilde{{\mathcal{G}}}_{q}$), recall of GPT-4o-identified relevant triples, and recall of answer entities within retrieved subgraphs or triples. The first metric assesses the ability of the approaches to retrieve the heuristic signals used as weak supervision during training. To provide a more accurate assessment for relevant triple retrieval, we employ GPT-4o to identify up to $20$ high-quality relevant triples for all test set samples, potentially capturing relevant triples beyond the shortest paths (see Appendix D for more details on labeling). We calculate individual recall values per question and average across all questions. For efficiency, we measure wall clock time on a 48GB NVIDIA RTX 6000 Ada GPU, reporting results in seconds. To focus on model computational efficiency, we exclude KG query times from our measurements. For text embedding computation, we only consider the time taken for question embedding computation as the entity and relation embeddings for the entire KG can be pre-computed.
@@
Overall Evaluation (Q1). Table 1 shows that SubgraphRAG consistently outperforms all other approaches in retrieval effectiveness. RoG achieves the most competitive baseline performance for retrieving shortest path triples and answer entities. However, when evaluated on GPT-4o-labeled triples, RoG’s performance drops significantly ($45.6\%$ for WebQSP, $52.2\%$ for CWQ), while SubgraphRAG remains robust ($2.0\%$ decrease for WebQSP, $3.6\%$ increase for CWQ). This stark difference, despite using the same training signals, empirically validates that SubgraphRAG’s individual triple selection mechanism allows for more flexible and effective subgraph extraction compared to RoG’s constrained path search approach. Regarding efficiency, SubgraphRAG is only slightly slower than the cosine similarity baseline on WebQSP while being one to two orders of magnitude faster than other baselines. For the cosine similarity baseline and SubgraphRAG, we report recall metrics based on the top-$100$ retrieved triples ($2.3\%$ of total candidate triples on average). This budget consistently yields robust reasoning performance across LLMs in subsequent experiments.
@@
Ablation Study for Design Options and Retrieval Size. To evaluate individual component contributions in SubgraphRAG, we conduct an ablation study with several variants. MLP is a structure-free variant employing only text embeddings. Given the prevalence of GNNs, we consider GraphSAGE (Hamilton et al., 2017), a popular GNN, to update entity representations prior to the MLP-based triple scoring. We further augment both MLP and GraphSAGE with a one-hot-encoding topic entity indicator (MLP + topic entity and GraphSAGE + topic entity). Following Sun et al. (2018), we incorporate Personalized PageRank (PPR) (Haveliwala, 2002), seeded from the topic entities, to integrate structural information (MLP + topic entity + PPR). To account for both the varying capabilities of downstream LLMs and inference cost/latency constraints, we evaluate these variants across a broad spectrum of retrieval sizes ($K$).
@@
Fig. 3 presents the results on CWQ, with baselines included for reference. Larger retrieval sizes uniformly improve recall across all variants. Equipped with DDE, SubgraphRAG outperforms other variants, even at the relatively small average retrieval sizes of the baselines. This demonstrates that SubgraphRAG’s superiority is not solely attributable to larger retrieval sizes. Regarding design options, the topic entity indicator invariably leads to an improvement. In contrast, GNN variants often result in performance degradation compared to their MLP counterparts. We suspect that the diffusion of semantic information introduces noise in triple selection. Finally, PPR fails to reliably yield improvements. For the results on WebQSP, see Appendix B.3, which are also consistent.
@@
Multi-Hop and Multi-Topic Questions (Q2). To evaluate the effectiveness of various approaches in capturing structural information for complex multi-hop and multi-topic questions, we group questions based on the number of hops and topic entities. Table 2 presents the performance breakdown by hop count. SubgraphRAG consistently outperforms other methods on WebQSP and achieves the best overall performance on CWQ. Notably, while the cosine similarity baseline and RoG demonstrate competitive performance for single-hop questions, their performance degrades significantly for multi-hop questions. Appendix B.4 provides a performance breakdown for single-topic and multi-topic questions, focusing exclusively on CWQ due to the predominance of single-topic questions in the WebQSP test set ($98.3\%$). SubgraphRAG consistently exhibits superior performance across all metrics for both single-topic and multi-topic questions. Our comprehensive analysis highlights the remarkable effectiveness of DDE in capturing complex topic-centered structural information essential for challenging questions involving multi-hop reasoning and multiple topic entities.

### v3

@@ 3.1 Efficient, Flexible and Expressive Subgraph Retrieval
A Lightweight Implementation For $p_{\theta}(\cdot|z_{\tau}({\mathcal{G}},q),q)$ We present a lightweight implementation of $p_{\theta}$ that integrates structural and semantic information. Following previous approaches (Karpukhin et al., 2020; Gao et al., 2024b), we employ off-the-shelf pre-trained text encoders to embed all entities/relations in a KG based on their text attributes. These semantic text embeddings are computed and stored in a vector database during the pre-processing stage for efficient retrieval. For a newly arrived question $q$, we embed $q$ to obtain $z_{q}$ and retrieve embeddings $z_{h},z_{r},z_{t}$ from the vector database for the involved entities and relations. After computing DDEs $z_{\tau}$, an MLP is employed for binary classification using the concatenated input $[z_{q}\|z_{h}\|z_{r}\|z_{t}\|z_{\tau}]$.
@@
Relevant Designs We considered several alternatives but found them less preferable due to concerns about efficiency and adaptability to KG updates. Cross-encoders, which concatenate a question and a retrieval candidate for joint embedding (Wolf et al., 2019), potentially offer better retrieval performance. However, due to the inability to pre-compute embeddings, this approach significantly reduces retrieval efficiency when dealing with a large number of retrieval candidates, as is the case in triple retrieval. Li et al. (2023a) embeds each triple as a whole rather than individual entities and relations. However, this approach incurs higher computational and storage costs and exhibits reduced generalizability to the triples that are new combinations of old entities and relations. Our lightweight implementation allows for fast triple scoring while maintaining good generalizability.
@@
Figure 2: The prompt used in SubgraphRAG. Concrete examples can be found in Appendix E).
@@ 3.2 Prompting-Based LLM Reasoning
We employ an LLM to reason over ${\mathcal{G}}_{q}$ by incorporating the linearized list of triples forming ${\mathcal{G}}_{q}$ into the prompt. This approach allows the LLM to ground its reasoning in the retrieved structured knowledge and select the answers $\hat{\mathcal{A}}_{q}$ from the entities within ${\mathcal{G}}_{q}$, mitigating issues like hallucinations and outdated knowledge (Lin et al., 2019; Shuster et al., 2021; Vu et al., 2024). Specifically, we prompt the LLM to generate knowledge-grounded explanations besides the answers, which further effectively reduces hallucinations. We adopt in-context learning (ICL) (Brown et al., 2020) and design dedicated prompt templates with explanation demonstrations to guide LLM reasoning (see Fig. 2).
Without fine-tuning LLMs, we reduce computational costs, enable the usage of state-of-the-art (SOTA) black-box LLMs, and maintain the generalizability for unseen KGs. Fine-tuning also risks enhancing prediction accuracy at the cost of general reasoning and explaining capabilities. High-quality text-based explanations are typically unavailable for question-answering tasks in practice. As a result, previous KG-based RAG approaches that rely on fine-tuning often depend on larger, unfine-tuned LLMs like GPT-4 to generate explanations as additional training data (Luo et al., 2024b).
Regarding the size $K$ of the retrieved subgraph, while increasing $K$ in principle improves the coverage of relevant information, it also risks introducing more irrelevant information that may harm LLM reasoning (Xu et al., 2024; Liu et al., 2024b). Different LLMs are inherently equipped with different sized context windows and also exhibit distinct capabilities in reasoning over long-context retrieval results (Dubey et al., 2024). As such, although the training of SubgraphRAG retriever is LLM-agnostic, the size $K$ needs to be properly selected per LLM and constraints. In Section 4.2, we empirically verify that more powerful LLMs can benefit from incorporating a larger-sized retrieved subgraph, demonstrating the benefit of size-adjustable subgraph retrieval in SubgraphRAG.
@@ 4 Experiments
Model Shortest Path Triple Recall GPT-4o Triple Recall Answer Entity Recall WebQSP CWQ WebQSP CWQ WebQSP CWQ $1$ $2$ $1$ $2$ $\geq 3$ $1$ $2$ $1$ $2$ $\geq 3$ $1$ $2$ $1$ $2$ $\geq 3$ $(65.8\%)$ $(34.2\%)$ $(28.0\%)$ $(65.9\%)$ $(6.1\%)$ $(65.8\%)$ $(34.2\%)$ $(28.0\%)$ $(65.9\%)$ $(6.1\%)$ $(65.8\%)$ $(34.2\%)$ $(28.0\%)$ $(65.9\%)$ $(6.1\%)$ cosine similarity $0.874$ $0.405$ $0.629$ $0.442$ $0.333$ $0.847$ $0.483$ $0.629$ $0.511$ $0.464$ $0.943$ $0.253$ $0.903$ $0.472$ $0.289$ SR+NSM w E2E $0.565$ $0.324$ $0.580$ $0.376$ $0.916$ $0.301$ Retrieve-Rewrite-Answer $0.064$ $0.046$ $0.062$ $0.061$ $0.745$ $0.729$ RoG $0.869$ $0.415$ $0.766$ $0.597$ $0.253$ $0.446$ $0.271$ $0.347$ $0.293$ $0.122$ $0.874$ $0.677$ $0.920$ $0.827$ $0.628$ G-Retriever $0.335$ $0.216$ $0.134$ $0.205$ $0.168$ $0.345$ $0.284$ $0.159$ $0.240$ $0.226$ $0.596$ $0.446$ $0.377$ $0.384$ $0.269$ GNN-RAG $0.532$ $0.502$ $0.515$ $0.498$ $0.446$ $0.384$ $0.445$ $0.328$ $0.408$ $0.418$ $0.810$ $0.831$ $0.853$ $0.841$ $0.787$ MLP $0.828$ $0.687$ $0.651$ $0.690$ $0.534$ $0.811$ $0.781$ $0.635$ $0.707$ $0.616$ $0.933$ $0.874$ $0.932$ $0.870$ $\underline{\mathbf{0.793}}$ MLP + topic entity $0.944$ $0.729$ $\underline{\mathbf{0.854}}$ $0.750$ $0.560$ $0.884$ $0.775$ $0.769$ $0.773$ $0.647$ $0.976$ $0.843$ $\underline{\mathbf{0.956}}$ $0.885$ $0.665$ SubgraphRAG $\underline{\mathbf{0.953}}$ $\underline{\mathbf{0.748}}$ $0.831$ $\underline{\mathbf{0.820}}$ $\underline{\mathbf{0.626}}$ $\underline{\mathbf{0.908}}$ $\underline{\mathbf{0.809}}$ $\underline{\mathbf{0.823}}$ $\underline{\mathbf{0.860}}$ $\underline{\mathbf{0.755}}$ $\underline{\mathbf{0.977}}$ $\underline{\mathbf{0.881}}$ $0.946$ $\underline{\mathbf{0.916}}$ $0.741$
@@ 4.1 Evaluation for Retrieval (Q1 & Q2)
Baseline Retrievers. Cosine similarity performs structure-free triple retrieval based on triple embeddings (Li et al., 2023a). SR+NSM w E2E (Zhang et al., 2022) and Retrieve-Rewrite-Answer (Wu et al., 2023) propose constrained path search based on pre-trained text encoders. RoG (Luo et al., 2024b) adopts a similar strategy but fine-tunes an LLM for relation path prediction. G-Retriever (He et al., 2024) combines cosine similarity search with combinatorial optimization to construct a connected subgraph. GNN-RAG employs lightweight GNNs to predict answer entities and extracts shortest paths between them and topic entities (Mavromatis & Karypis, 2022; 2024).
Implementation Details. We adopt gte-large-en-v1.5 (Li et al., 2023c) as the pre-trained text encoder for both the cosine similarity baseline and SubgraphRAG, a 434M model that achieves a good balance between efficiency and English retrieval performance, as evidenced by the Massive Text Embedding Benchmark (MTEB) leaderboard (Muennighoff et al., 2023). For supervision signals, there are no ground-truth relevant subgraphs for a query $q$. Previous path-based subgraph retrievers adopt the shortest paths between the topic and answer entities for the weak supervision signals (Zhang et al., 2022; Luo et al., 2024b). We also utilize these shortest paths as the heuristic relevant subgraphs $\tilde{{\mathcal{G}}}_{q}$ to train $\mathbb{Q}_{\theta}$ as in Eq. 2. To reduce the size of candidate triples ${\mathcal{G}}$, we construct subgraphs centered at topic entities ${\mathcal{T}}_{q}$, following previous works. See Appendix C.1 for more details.
@@
WebQSP CWQ Macro-F1 Hit Macro-F1 Hit UniKGQA $72.2$ $49.0$ SR+NSM w E2E $64.1$ $46.3$ KD-CoT 52.5 68.6 - 55.7 ToG (GPT-4) - 82.6 - 67.6 StructGPT - 74.69 - - Retrieve-Rewrite-Answer - 79.36 - - G-Retriever 53.41 73.46 - - RoG-Joint 70.26 86.67 54.63 61.94 RoG-Sep 66.45 82.19 53.87 60.55 EtD - $82.5$ $62.0$ GNN-RAG $71.3$ $85.7$ $59.4$ $66.8$ SubgraphRAG + Llama3.1-8B 70.57 86.61 47.16 56.98 SubgraphRAG + Llama3.1-70B 74.70 86.24 51.78 57.89 SubgraphRAG + ChatGPT 69.21 83.11 49.13 56.27 SubgraphRAG + GPT4o-mini 77.45 90.11 54.13 62.02 SubgraphRAG + GPT4o 76.46 89.80 59.08 66.69 SubgraphRAG + GPT4o-mini (200) 77.82 90.54 54.69 63.49 SubgraphRAG + GPT4o (200) 78.24 90.91 59.42 67.49 SubgraphRAG + GPT4o-mini (500) 77.67 91.22 55.41 64.97 SubgraphRAG + Llama3.1-8B ($\leftrightarrow$) 66.42 83.42 37.96 48.57 SubgraphRAG + GPT4o-mini ($\leftrightarrow$) 73.81 88.08 44.69 54.21 SubgraphRAG + GPT4o-mini ($\leftrightarrow$, 500) 76.20 91.22 50.30 60.80 SubgraphRAG + GPT4o ($\leftrightarrow$) 74.27 87.10 47.55 54.89 SubgraphRAG + GPT4o ($\leftrightarrow$, 500) 77.12 89.68 55.14 62.98
@@
Evaluation Metrics. Our evaluation covers both effectiveness and efficiency. For effectiveness, we employ three recall metrics, respectively for shortest path triples (i.e., $\tilde{{\mathcal{G}}}_{q}$), GPT-4o-identified relevant triples, and answer entities within retrieved subgraphs. The first metric assesses the retrieval of weak supervision signals used for training. For a more accurate assessment, we employ GPT-4o to identify $\leq 20$ relevant triples for all test samples, potentially capturing relevant triples beyond the shortest paths (see Appendix E for labeling details). We calculate individual recall values per question and then take the average. For efficiency, we measure wall clock time on a 48GB NVIDIA RTX 6000 Ada GPU in seconds. To focus on model computational efficiency, we exclude KG query time. For text embedding, employed by SubgraphRAG variants and most baselines (cosine similarity, G-Retriever, and GNN-RAG), we only count the time for question embedding as the entity and relation embeddings for the entire KG can be pre-computed. Such pre-computation may yield a time-memory trade-off.
@@
Overall Evaluation (Q1). Table 1 shows that SubgraphRAG consistently outperforms all other approaches in retrieval effectiveness. RoG achieves the most competitive baseline performance for retrieving shortest path triples. However, when evaluated on GPT-4o-labeled triples, RoG’s performance drops significantly ($45.6\%$ for WebQSP, $52.2\%$ for CWQ), while SubgraphRAG remains robust ($2.0\%$ decrease for WebQSP, $3.6\%$ increase for CWQ). This stark difference, despite using the same training signals, empirically validates that SubgraphRAG’s individual triple selection mechanism allows for more flexible and effective subgraph extraction compared to RoG’s constrained path search approach. Regarding efficiency, SubgraphRAG is only slightly slower than the cosine similarity baseline on WebQSP while being one to two orders of magnitude faster than other baselines, including GNN-RAG. For the cosine similarity baseline and SubgraphRAG, we report recall metrics based on the top-$100$ retrieved triples ($2.3\%$ of total candidate triples on average). This budget consistently yields robust reasoning performance across LLMs in subsequent experiments.
@@
Ablation Study for Design Options and Retrieval Size. To evaluate individual component contributions, we conduct an ablation study with several variants. MLP is a structure-free variant employing only text embeddings. We consider GraphSAGE (Hamilton et al., 2017), a popular GNN, to update entity representations prior to the triple scoring. We further augment both MLP and GraphSAGE with a one-hot-encoding topic entity indicator (MLP + topic entity and GraphSAGE + topic entity). Following Sun et al. (2018), we incorporate Personalized PageRank (PPR) (Haveliwala, 2002), seeded from the topic entities, to integrate structural information (MLP + topic entity + PPR). To account for both the varying capabilities of downstream LLMs and inference cost/latency constraints, we evaluate these variants across a broad spectrum of retrieval sizes ($K$).
@@
Fig. 3 presents the results on CWQ, with baselines included for reference. Larger retrieval sizes uniformly improve recall across all variants. Equipped with DDE, SubgraphRAG outperforms other variants, even at the relatively small average retrieval sizes of the baselines. This demonstrates that SubgraphRAG’s superiority is not solely attributable to larger retrieval sizes. Regarding design options, the topic entity indicator invariably leads to an improvement. In contrast, GNN variants often result in performance degradation compared to their MLP counterparts. We suspect that the diffusion of semantic information introduces noise in triple selection. Finally, PPR fails to reliably yield improvements. For the results on WebQSP, see Appendix C.3, which are also consistent.
@@
Multi-Hop and Multi-Topic Questions (Q2). To evaluate the effectiveness of various approaches in capturing structural information for complex multi-hop and multi-topic questions, we group questions based on the number of hops and topic entities. Table 2 presents the performance breakdown by hop count. SubgraphRAG consistently outperforms other methods on WebQSP and achieves the best overall performance on CWQ. Notably, while the cosine similarity baseline and RoG demonstrate competitive performance for single-hop questions, their performance degrades significantly for multi-hop questions. Appendix C.4 provides a performance breakdown for single-topic and multi-topic questions, focusing exclusively on CWQ due to the predominance of single-topic questions in the WebQSP test set ($98.3\%$). SubgraphRAG consistently exhibits superior performance across all metrics for both single-topic and multi-topic questions. Our comprehensive analysis highlights the remarkable effectiveness of DDE in capturing complex topic-centered structural information essential for challenging questions involving multi-hop reasoning and multiple topic entities.
4.2 KGQA Results (Q3 & Q4)
KGQA Baselines. Besides the baselines in retrieval evaluation, we include more KGQA methods, including UniKGQA (Jiang et al., 2023b), KD-CoT (Wang et al., 2023a), StructGPT (Jiang et al., 2023a), ToG (Sun et al., 2024a), and EtD (Liu et al., 2024a). For RoG, we present two entries: RoG-Joint and RoG-Sep. Originally, RoG fine-tuned its LLMs on the training sets of both WebQSP and CWQ. Yet, this joint training approach leads to significant label leakage, with over 50% of WebQSP test questions (or their variants) appearing in CWQ’s training set, and vice versa. Therefore, we re-trained RoG on each dataset separately, denoted by RoG-Sep.
Evaluation Metrics. Along with the commonly reported Macro-F1 and Hit [^1] The baselines claim to report Hit@1, but they actually compute Hit, which measures whether at least one correct answer appears in the LLM response., we also include Micro-F1 to account for the imbalance in the number of ground-truth answers across samples and Hit@1 for a more inclusive evaluation. To further assess model performance, we introduce scoreh, inspired by (Yang et al., 2024), which evaluates how truth-grounded the predicted answers are and the degree of hallucination. This metric penalizes hallucinated answers while favoring missing answers over incorrect ones. Scores are normalized to a range of 0 to 100, with higher scores indicating better truth-grounding in the model’s answers. The scoring strategy is described in detail in Appendix D.